# Complete Characterizations of Well-Posedness in Parametric Composite Optimization

### v1

@@ Abstract
This paper provides complete characterization of well-posedness for Karush–Kuhn–Tucker (KKT) systems associated with general problems of perturbed composite optimization. Leveraging the property of parabolic regularity for composite models, we show that the second-order subderivative of the cost function reduces to the novel second-order variational function playing a crucial role in the subsequent analysis. This foundational result implies that the strong second-order sufficient condition (SSOSC) introduced in this work for the general class of composite optimization problems naturally extends the classical second-order sufficient condition in nonlinear programming. Then we obtain several equivalent characterizations of the second-order qualification condition (SOQC) and highlight its equivalence to the constraint nondegeneracy condition under the $\mathcal{C}^{2}$-cone reducibility assumption. These insights lead us to multiple equivalent conditions for the major Lipschitz-like/Aubin property of KKT systems, including the SOQC combined with the new second-order subdifferential condition and the SOQC combined with tilt stability of local minimizers. Furthermore, under $\mathcal{C}^{2}$-cone reducibility, we prove that the Lipschitz-like property of the reference KKT system is equivalent to its strong regularity. Finally, we demonstrate that the Lipschitz-like property is equivalent to the nonsingularity of the generalized Jacobian associated with the KKT system under a certain verifiable assumption. These results provide a unified and rigorous framework for analyzing stability and sensitivity of solutions to composite optimization problems, as well as for the design and justification of numerical algorithms.
@@ 3 Qualification and Second-Order Variational Conditions
It follows from [Bonnans2000, Theorem 3.9 and Proposition 3.17] that the Lagrange multiplier set $\Lambda(x,0,0)$ is nonempty, convex, and compact if and only if the RCQ holds at $x$. If $\bar{x}$ is a local minimizer of problem (1.1) and the RCQ is satisfied at $\bar{x}$, then there exists a Lagrange multiplier $\bar{u}\in\mathcal{R}^{m}$ such that $(\bar{x},\bar{u})\in S_{\operatorname{KKT}}(0,0)$. It is said that the strong Robinson constraint qualification (SRCQ) is satisfied at $\bar{x}$ with respect to $\bar{u}$ if we have
@@ 5 Characterizations of Lipschitzian Stability for KKT Systems
Let $(x,u)$ solve the KKT system (1.3) with parameters $(a,0)$. Suppose that $g:\mathcal{R}^{m}\to\overline{\mathcal{R}}$ satisfies assumption (A1) at $F(x)$. Then the mapping $S_{\operatorname{KKT}}$ has the isolated calmness property at $(a,0)$ for $(x,u)$ if and only if both SRCQ (3.3) and the following version
@@
$\left[\begin{array}[]{c}0\\ 0\end{array}\right]$ $\in\left[\begin{array}[]{c}\nabla_{xx}^{2}L(x,u)\Delta\xi+\nabla F(x)^{T}\Delta\eta\\ -\Delta\eta+D(\partial g)(F(x),u)(\nabla F(x)\Delta\xi)\end{array}\right]\Longleftrightarrow$
$\left[\begin{array}[]{c}0\\ 0\end{array}\right]\in\left[\begin{array}[]{c}\nabla_{xx}^{2}L(x,u)\Delta\xi+\nabla F(x)^{T}\Delta\eta\\ -\nabla F(x)\Delta\xi+D(\partial g^{*})(u,F(x))(\Delta\eta)\end{array}\right].$
It follows from Lemma 2.2 that the solution mapping $S_{\operatorname{KKT}}$ enjoys the isolated calmness property at $(a,0)$ for $(x,u)$ if and only if the modified KKT mapping $\widetilde{S}_{\operatorname{KKT}}$ has this property at $((0,0),(0,0))$ for ((x,c),(u,1))((x,c),(u,-1)). By the above expression of $d^{2}\delta_{\operatorname{epi}g}$, the original SOSC (3.6) for (5.2) is equivalent to the SOSC version in (5.1). The equivalence between the SRCQ for (5.2) and the original SRCQ (3.3) is trivial. The proof is completed by applying [MohammadiMordukhovichSarabi2020, Theorem 4.1]. ∎
@@
Let $(\bar{x},\bar{u})$ solve the KKT system (1.3) with $(a,b)=(0,0)$. Suppose that RCQ (3.1) holds at $\bar{x}$, that $g$ satisfies (A1) at $F(\bar{x})$, and that there exist neighborhoods $\mathcal{U}$ of $\bar{x}$ and $\mathcal{V}$ of the origin such that $\Lambda(x,a,0)$ is a singleton for $x\in\mathcal{U}$, $a\in\mathcal{V}$, and $(x,u)\in S_{\operatorname{KKT}}(a,0)$. If $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ with modulus $\kappa>0$, then we have
@@
Let $(\bar{x},\bar{u})$ be a solution to the KKT system (1.3) with $(a,b)=(0,0)$. Suppose that SOQC (4.1) holds at $\bar{x}$ for $\bar{u}$. Then there exist neighborhoods $\mathcal{U}$ of $\bar{x}$ and $\mathcal{V}$ of the origin such that the SOQC is satisfied at $x$ for $u$ whenever $x\in\mathcal{U}$, $a\in\mathcal{V}$ with $(x,u)\in S_{\operatorname{KKT}}(a,0)$.
@@
Assuming the contrary, take any sequence $a^{k}\rightarrow 0$ and find $(x^{k},u^{k})\in S_{\operatorname{KKT}}(a^{k},0)$ with $x^{k}\rightarrow\bar{x}$ and $w^{k}\neq 0$ with $\|w^{k}\|=1$ such that
@@
If $(A1)$ is fulfilled for $g$ at $F(\bar{x})$, the above are equivalent to each of the properties:
@@
$\begin{array}[]{ll}D^{*}\Phi((x,u),(\delta_{1},0))(\Delta\xi,\Delta\eta)&=\left[\begin{array}[]{c}\nabla^{2}_{xx}L(x,u)\Delta\xi-\nabla F(x)^{T}\Delta\eta\\ \nabla F(x)\Delta\xi\end{array}\right]\\ \\ &+\left[\begin{array}[]{c}0\\ I\end{array}\right]D^{*}(\partial g^{*})(u,F(x))(\Delta\eta).\end{array}$ (5.14)
By the Mordukhovich criterion in Lemma 2.1, the Aubin property of $S_{\operatorname{KKT}}$ at $(\delta_{1},0)$ for $(x,u)$ is equivalent to the implication $(0,0)\in D^{*}\Phi((x,u),(\delta_{1},0))(\Delta\xi,\Delta\eta)\Longrightarrow(\Delta\xi,\Delta\eta)=(0,0)$. This tells us by (5.14) that if $\Delta\zeta\in D^{*}(\partial g^{*})(u,F(x))(\Delta\eta)$ with
@@
$0\in D^{*}\Phi((x,u),(\delta_{1},0))(0,\Delta\eta)\ \Longleftrightarrow\ \nabla F(x)^{T}\Delta\eta=0,\ -\Delta\eta\in D^{*}(\partial g)(F(x),u)(0)$
@@
$D\Phi((x,u),(\delta_{1},0))(\Delta\xi^{\prime},\Delta\eta^{\prime})=\left[\begin{array}[]{c}\nabla_{xx}^{2}L(x,u)\Delta\xi^{\prime}+\nabla F(x)^{T}\Delta\eta^{\prime}\\ -\nabla F(x)\Delta\xi^{\prime}+D(\partial g^{*})(u,F(x))(\Delta\eta^{\prime})\end{array}\right].$
Hence the inclusion $(0,0)\in D\Phi((x,u),(\delta_{1},0))(\Delta\xi^{\prime},\Delta\eta^{\prime})$ yields $\nabla_{xx}^{2}L(x,u)\Delta\xi^{\prime}+\nabla F(x)^{T}\Delta\eta^{\prime}=0$. It follows from [MordukhovichTangWang-tilt-2025, Proposition 3.2] that $g$ is twice epi-differentiable at the points in question. This leads us by using (2.11), (2.12), and [Rockafellar1998, Theorem 13.57] that
@@
and thus $(0,0)\in D^{*}\Phi((x,u),(\delta_{1},0))(\Delta\xi^{\prime},-\Delta\eta^{\prime})$. Employing the characterization of Lemma 2.1, we deduce from the Aubin property of $S_{\operatorname{KKT}}$ at $(\delta_{1},0)$ for $(x,u)$ that $\Delta\xi^{\prime}=0$ and $\Delta\eta^{\prime}=0$. Then the Levy-Rockafellar criterion in Lemma 2.2 tells us that $S_{\operatorname{KKT}}$ enjoys the isolated calmness property at $(\delta_{1},0)$ for $(x,u)$. Lemma 5.1, valid under (A1), confirms that SRCQ (3.3) and SOSC (5.1) hold for $(x,u)\in\Phi^{-1}(\delta_{1},0)$ with $\delta_{1}\in\mathcal{V}$ and $(x,u)\in\mathcal{U}$. Deduce from [MordukhovichTangWang-tilt-2025, Theorem 4.1] the existence of $\kappa>0$ such that $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ with modulus $\kappa$, which justifies assertion (iv).
@@
Let $\bar{x}$ be a local minimizer of problem (1.1), where $g$ is $\mathcal{C}^{2}$-cone reducible and the Jacobian $\nabla F(\bar{x})$ is of full rank. Then conditions (5.6) and (5.17) are equivalent. In the case where $F(x):=x$, we have that the Aubin property of the generalized equation
@@
at the origin for $\bar{x}$ is equivalent to the strong regularity of $\bar{x}$ for (5.27).
@@
Let $(\bar{x},\bar{u})$ be a solution to the KKT system (1.3) with $(a,b)=(0,0)$, where $g$ satisfies (A1) and (A2) at $F(\bar{x})$ for $\bar{u}$. If SOQC (4.1) and SSOSC (3.10) hold at $\bar{x}$, then all the matrices from the generalized Jacobian $\mathcal{J}Z(\bar{x},\bar{u})$ are nonsingular.
@@
Let $(\bar{x},\bar{u})$ solve the KKT system (1.3) with $(a,b)=(0,0)$, and let $g$ satisfy both assumptions (A1), (A2) at $F(\bar{x})$ for $\bar{u}$. Then we have the equivalent assertions:
@@
(ii) SOQC (4.1) holds together with the SSOSC at $(\bar{x},\bar{u})$ given by
@@
Let $(\bar{x},\bar{u})$ solve the KKT system (1.3) with $(a,b)=(0,0)$, and let $g$ be $\mathcal{C}^{2}$-cone reducible while satisfying (A2) at $F(\bar{x})$ for $\bar{u}$. Then the following assertions are equivalent:

### v2

@@ Abstract
This paper provides complete characterization of well-posedness for Karush–Kuhn–Tucker (KKT) systems associated with general problems of perturbed composite optimization. Leveraging the property of parabolic regularity for composite models, we show that the second-order subderivative of the cost function reduces to the novel second-order variational function playing a crucial role in the subsequent analysis. This foundational result implies that the strong second-order sufficient condition introduced in this work for the general class of composite optimization problems naturally extends the classical second-order sufficient condition in nonlinear programming. Then we obtain several equivalent characterizations of the second-order qualification condition (SOQC) and highlight its equivalence to the constraint nondegeneracy condition under the $\mathcal{C}^{2}$-cone reducibility assumption. These insights lead us to multiple equivalent conditions for the major Lipschitz-like/Aubin property of KKT systems, including SOQC combined with the new second-order subdifferential condition and SOQC combined with tilt stability of local minimizers. Furthermore, under $\mathcal{C}^{2}$-cone reducibility, we prove that the Lipschitz-like property of the reference KKT system is equivalent to its strong regularity. Finally, we demonstrate that the Lipschitz-like property is equivalent to the nonsingularity of the generalized Jacobian associated with the KKT system under a certain verifiable assumption. These results provide a unified and rigorous framework for analyzing stability and sensitivity of solutions to composite optimization problems, as well as for the design and justification of numerical algorithms.
@@ 3 Qualification and Second-Order Variational Conditions
It follows from [Bonnans2000, Theorem 3.9 and Proposition 3.17] that the Lagrange multiplier set $\Lambda(x,0,0)$ is nonempty, convex, and compact if and only if RCQ holds at $x$. If $\bar{x}$ is a local minimizer of problem (1.1) and RCQ is satisfied at $\bar{x}$, then there exists a Lagrange multiplier $\bar{u}\in\mathcal{R}^{m}$ such that $(\bar{x},\bar{u})\in S_{\operatorname{KKT}}(0,0)$. It is said that the strong Robinson constraint qualification (SRCQ) is satisfied at $\bar{x}$ with respect to $\bar{u}$ if we have
@@ 5 Characterizations of Lipschitzian Stability for KKT Systems
Let $x$ be a local minimizer of problem (1.2) with parameters $(a,0)$ and $u\in\Lambda(x,a,0)$ be a corresponding Lagrange multiplier. Suppose that $g:\mathcal{R}^{m}\to\overline{\mathcal{R}}$ satisfies assumption (A1) at $F(x)$. Then the mapping $S_{\operatorname{KKT}}$ has the isolated calmness property at $(a,0)$ for $(x,u)$ if and only if both SRCQ (3.3) and the following version
@@
$\left[\begin{array}[]{c}0\\ 0\end{array}\right]$ $\in\left[\begin{array}[]{c}\nabla_{xx}^{2}L(x,u)\Delta\xi+\nabla F(x)^{T}\Delta\eta\\ -\Delta\eta+D(\partial g)(F(x),u)(\nabla F(x)\Delta\xi)\end{array}\right]\Longleftrightarrow\left[\begin{array}[]{c}0\\ 0\end{array}\right]\in\left[\begin{array}[]{c}\nabla_{xx}^{2}L(x,u)\Delta\xi+\nabla F(x)^{T}\Delta\eta\\ -\nabla F(x)\Delta\xi+D(\partial g^{*})(u,F(x))(\Delta\eta)\end{array}\right].$
It follows from Lemma 2.2 that the solution mapping $S_{\operatorname{KKT}}$ enjoys the isolated calmness property at $(a,0)$ for $(x,u)$ if and only if the modified KKT mapping $\widetilde{S}_{\operatorname{KKT}}$ has this property at $((0,0),(0,0))$ for ((x,c),(u,1))((x,c),(u,-1)). By the above expression of $d^{2}\delta_{\operatorname{epi}g}$, the original SOSC (3.6) for (5.2) is equivalent to the SOSC version in (5.1). The equivalence between SRCQ for (5.2) and the original SRCQ (3.3) is trivial. Due to the fact that $(x,g(F(x)))$ is also the unique local minimizer of problem (5.2), the proof is completed by applying [MohammadiMordukhovichSarabi2020, Theorem 4.1]. ∎
@@
Let $\bar{x}$ be a local minimizer of problem (1.1) with $\bar{u}\in\Lambda(\bar{x},0,0)$ being a corresponding Lagrange multiplier. Suppose that RCQ (3.1) holds at $\bar{x}$, that $g$ satisfies (A1) at $F(\bar{x})$, and that there exist neighborhoods $\mathcal{U}$ of $\bar{x}$ and $\mathcal{V}$ of the origin such that $\Lambda(x,a,0)$ is a singleton for $x\in\mathcal{U}$, $a\in\mathcal{V}$, and $(x,u)\in S_{\operatorname{KKT}}(a,0)$. If $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ with modulus $\kappa>0$, then we have
@@
Let $(\bar{x},\bar{u})$ be a solution to the KKT system (1.3) with $(a,b)=(0,0)$. Suppose that SOQC (4.1) holds at $\bar{x}$ for $\bar{u}$. Then there exist neighborhoods $\mathcal{U}$ of $\bar{x}$ and $\mathcal{V}$ of the origin such that SOQC is satisfied at $x$ for $u$ whenever $x\in\mathcal{U}$, $\delta\in\mathcal{V}$ with $(x,u)\in S_{\operatorname{KKT}}(\delta,0)$.
@@
Assuming the contrary, for any $k$ sufficiently take a sequence $\delta^{k}\in\frac{1}{k}\mathbb{B}_{\mathcal{R}^{n}}$ and find $(x^{k},u^{k})\in S_{\operatorname{KKT}}(\delta^{k},0)$ with $x^{k}\rightarrow\bar{x}$ and $w^{k}\neq 0$ with $\|w^{k}\|=1$ such that
@@
If $\bar{x}$ is a local minimizer of problem (1.1), $(A1)$ is fulfilled for $g$ at $F(\bar{x})$, the above are equivalent to each of the properties:
@@
$\begin{array}[]{ll}D^{*}\Phi((x,u),(\delta,0))(\Delta\xi,\Delta\eta)&=\left[\begin{array}[]{c}\nabla^{2}_{xx}L(x,u)\Delta\xi-\nabla F(x)^{T}\Delta\eta\\ \nabla F(x)\Delta\xi\end{array}\right]+\left[\begin{array}[]{c}0\\ I\end{array}\right]D^{*}(\partial g^{*})(u,F(x))(\Delta\eta).\end{array}$ (5.7)
By the Mordukhovich criterion in Lemma 2.1, the Aubin property of $S_{\operatorname{KKT}}$ at $(\delta,0)$ for $(x,u)$ is equivalent to the implication $(0,0)\in D^{*}\Phi((x,u),(\delta,0))(\Delta\xi,\Delta\eta)\Longrightarrow(\Delta\xi,\Delta\eta)=(0,0)$. This tells us by (5.7) that if $\Delta\zeta\in D^{*}(\partial g^{*})(u,F(x))(\Delta\eta)$ with
@@
$0\in D^{*}\Phi((x,u),(\delta,0))(0,\Delta\eta)\ \Longleftrightarrow\ \nabla F(x)^{T}\Delta\eta=0,\ -\Delta\eta\in D^{*}(\partial g)(F(x),u)(0)$
@@
$0\in\nabla_{xx}^{2}L(x^{k},u^{k})d^{k}+\nabla F(x^{k})^{T}D^{*}(\partial g)(F(x^{k}),u^{k})(\nabla F(x^{k})d^{k}).$
By the outer semicontinuous of the coderivative operator, passing to the limit in the above inclusion and assuming without loss of generality $d^{k}\rightarrow\bar{d}$ with $\|\bar{d}\|=1$, it yields that
$0\in\nabla_{xx}^{2}L(\bar{x},\bar{u})\bar{d}+\nabla F(\bar{x})^{T}D^{*}(\partial g)(F(\bar{x}),\bar{u})(\nabla F(\bar{x})\bar{d}),$
[…]
$D\Phi((x,u),(\delta,0))(\Delta\xi^{\prime},\Delta\eta^{\prime})=\left[\begin{array}[]{c}\nabla_{xx}^{2}L(x,u)\Delta\xi^{\prime}+\nabla F(x)^{T}\Delta\eta^{\prime}\\ -\nabla F(x)\Delta\xi^{\prime}+D(\partial g^{*})(u,F(x))(\Delta\eta^{\prime})\end{array}\right].$
Hence the inclusion $(0,0)\in D\Phi((x,u),(\delta,0))(\Delta\xi^{\prime},\Delta\eta^{\prime})$ yields $\nabla_{xx}^{2}L(x,u)\Delta\xi^{\prime}+\nabla F(x)^{T}\Delta\eta^{\prime}=0$. It follows from [MordukhovichTangWang-tilt-2025, Proposition 3.2] that $g$ is twice epi-differentiable at the points in question. This leads us by using (2.11), (2.12), and [Rockafellar1998, Theorem 13.57] that
@@
and thus $(0,0)\in D^{*}\Phi((x,u),(\delta,0))(\Delta\xi^{\prime},-\Delta\eta^{\prime})$. Employing the characterization of Lemma 2.1, we deduce from the Aubin property of $S_{\operatorname{KKT}}$ at $(\delta,0)$ for $(x,u)$ that $\Delta\xi^{\prime}=0$ and $\Delta\eta^{\prime}=0$. Then the Levy-Rockafellar criterion in Lemma 2.2 tells us that $S_{\operatorname{KKT}}$ enjoys the isolated calmness property at $(\delta,0)$ for $(x,u)$ and $S_{\operatorname{KKT}}(\delta,0)$ is a singleton. Let $\delta=0$, Lemma 5.1, valid under (A1), confirms that SOQC (4.1) and SOSC (5.1) hold for $(\bar{x},\bar{u})$. Consequently, this guarantees that the second-order growth condition is satisfied for problem (1.1) at $\bar{x}$ (see, e.g., [Mohammadi2020, Theorem 6.1] ) and RCQ (3.1) holds at $\bar{x}$. It is straightforward to verify that the second-order growth condition for problem (4.7) also holds at the point $(\bar{x},g(F(\bar{x})))$. Applying [Bonnans2000, Proposition 4.32], we conclude that the solution set of problem (5.2) with the parameter $a=\delta$ is nonempty. This result implies that the solution set of problem (1.2) with parameters $(a,b)=(\delta,0)$ is also nonempty. Given any $\delta$ with its norm sufficiently small, let $x^{\delta}$ be a local minimizer of problem (1.2) with $(a,b)=(\delta,0)$. By [Bonnans2000, Proposition 4.32], $x^{\delta}$ is sufficiently close to $\bar{x}$ and RCQ holds at $x^{\delta}$. There exists a unique $u^{\delta}\in\Lambda(x^{\delta},\delta,0)$ with $\Lambda(x^{\delta},\delta,0)=\{u^{\delta}\}$ and $S_{\operatorname{KKT}}(\delta,0)=\Phi^{-1}(\delta,0)=\{(x^{\delta},u^{\delta})\}$ both singletons. We conclude that for any $(x,u)\in S_{\operatorname{KKT}}(\delta,0)$, $x$ is a local minimizer of problem (1.2) with $(a,b)=(\delta,0)$ and $S_{\operatorname{KKT}}$ has the isolated calmness property at $(\delta,0)$ for $(x,u)$. Applying Lemma 5.1, it follows that SOQC (4.1) and SOSC (5.1) hold for any $(x,u)\in\Phi^{-1}(\delta,0)=\{(x^{\delta},u^{\delta})\}$ with $\delta\in\mathcal{V}$ and $(x,u)\in\mathcal{U}$. Deduce from [MordukhovichTangWang-tilt-2025, Theorem 4.1] the existence of $\kappa>0$ such that $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ with modulus $\kappa$, which justifies assertion (iv).
@@
Let $\bar{x}$ be a local minimizer of problem (1.1), where $g$ is $\mathcal{C}^{2}$-cone reducible and the Jacobian $\nabla F(\bar{x})$ is of full rank. Then conditions (5.6) and (5.10) are equivalent. In the case where $F(x):=x$, we have that the Aubin property of the solution mapping to the generalized equation
@@
at the origin for $\bar{x}$ is equivalent to the strong regularity of $\bar{x}$ for $\Psi^{-1}$ of (5.20).
@@
Let $\bar{x}$ be a local minimizer of problem (1.1) with $\bar{u}\in\Lambda(\bar{x},0,0)$ as a corresponding Lagrange multiplier, where $g$ satisfies (A1) and (A2) at $F(\bar{x})$ for $\bar{u}$. If SOQC (4.1) and SSOSC (3.10) hold at $\bar{x}$, then all the matrices from the generalized Jacobian $\mathcal{J}Z(\bar{x},\bar{u})$ are nonsingular.
@@
Let $\bar{x}$ be a local minimizer of problem (1.1) with $\bar{u}\in\Lambda(\bar{x},0,0)$ as a corresponding Lagrange multiplier, and let $g$ satisfy both assumptions (A1), (A2) at $F(\bar{x})$ for $\bar{u}$. Then we have the equivalent assertions:
@@
(ii) SOQC (4.1) holds together with SSOSC at $(\bar{x},\bar{u})$ given by
@@
$\bar{x}$ be a local minimizer of problem (1.1) with $\bar{u}\in\Lambda(\bar{x},0,0)$ as a corresponding Lagrange multiplier, and let $g$ be $\mathcal{C}^{2}$-cone reducible while satisfying (A2) at $F(\bar{x})$ for $\bar{u}$. Then the following assertions are equivalent: